# A Survey on Unlearning in Large Language Models

### v1

@@ 4.1.3. Experiment Paradigm
In the experiment, datasets can be broadly categorized into two classes based on whether fine-tuning is required. In the first category, the models perform unlearning without fine-tuning on the target dataset, which simplifies experimental setup. This includes the following scenarios. (1) The unlearn set is compiled directly from the model’s pretraining data, such as subsets derived from the Pile (Gao et al., 2020). (2) The data are manually verified to be present in the model, as in RWKU (Jin et al., 2024), ConceptVectors (Hong et al., 2024), and RETURN (Liu et al., 2025b). However, verifying the presence of facts in LLMs remains challenging and may affect reliability. (3) For security purposes, the model is required to erase certain knowledge regardless of its original presence, exemplified by WMDP (Li et al., 2024) and UNCD (Lang et al., 2025). In the second category, models are first fine-tuned on the full dataset before a subset is unlearned. This is essential when datasets are fictionally synthetic, such as TOFU (Maini et al., 2024), EDU-RELAT (Wu et al., 2025), and PISTOL (Qiu et al., 2024), to ensure that the model acquires the target knowledge. Even for real-world corpora, fine-tuning helps to guarantee that the original model possesses knowledge of the unlearn set.
@@
TOFU (Maini et al., 2024) $0×\times2$0 QA SAQA (Joshi et al., 2024; Yuan et al., 2025b; Scholten et al., 2025; Wang et al., 2025b; Shen et al., 2025; Zhang et al., 2025a; Gao et al., 2025; Cha et al., 2024; Ren et al., 2025; Thaker et al., 2024; Huang et al., 2025; Liu et al., 2024a; Ji et al., 2024; Sanyal and Mandal, 2025; Takashiro et al., 2024; Jia et al., 2025; Xu et al., 2025; Gu et al., 2024; Ma et al., 2025; Wang et al., 2025d; Mekala et al., 2025; Jia et al., 2024; Wang et al., 2024; Chen et al., 2025; Liu et al., 2024b)
WMDP (Li et al., 2024) Papers & passages 3,668 MCQA (Che et al., 2025; Doshi and Stickland, 2024; Wang et al., 2025b; Farrell et al., 2024; Ren et al., 2025; Thaker et al., 2024; Liu et al., 2024a; Bhaila et al., 2025; Sanyal and Mandal, 2025; Kolbeinsson et al., 2025; Jia et al., 2025; Hu et al., 2025; Dang et al., 2025; Łucki et al., 2025)
WHP (Eldan and Russinovich, 2024) 3.1M tokens 300 Conti + 30 Cloze (Lynch et al., 2024; Scholten et al., 2025; Wang et al., 2025b; Thaker et al., 2024; Ji et al., 2024; Sanyal and Mandal, 2025; Jia et al., 2025, 2024; Chen et al., 2025)
MUSE (Shi et al., 2025) 4.4M+6.5M tokens Conti + SAQA (Jiang et al., 2025b; Wang et al., 2025b; Zhang et al., 2025e; Yuan et al., 2025a; Russinovich and Salem, 2025)
RWKU (Jin et al., 2024) 200 celebrities 3,268 cloze + 2,879 SAQA (Takashiro et al., 2024; Yuan et al., 2025a)
CoTaEval (Wei et al., 2025a) 1K + 1K passages 1.5K Conti + 1K SAQA (Russinovich and Salem, 2025)
KnowUnDo (Tian et al., 2024) 2,649 QA SAQA (Xu et al., 2025)
PISTOL (Qiu et al., 2024) 4 Graphs (50,95) 95 SAQA (Shen et al., 2025)
LUME (Ramakrishna et al., 2025) 1,387 documents 4,394 (Conti + SAQA) SEMEval-2025 Task 4
@@
ELUDe (Choi et al., 2025) 15,651+90,954 QA MCQA + SAQA -
FaithUn (Yang et al., 2025) 664 QA 8,377 MCQA -
LLM Surgery (Veldanda et al., 2024) 180K+1B tokens 24,800 MCQA -
Restor (Rezaei et al., 2024) 3,000 passages 1,051 SAQA -
@@
UNCD (Lang et al., 2025) 2.9M+3.3M tokens 36K MCQA -
WPU (Liu et al., 2024b) 100 people’s Wiki 2,795 SAQA -
(b) Outlines of benchmarks and datasets.
Table 4. Select a suitable benchmark. Part (a) organizes advantages of different tasks formats, data content and experiment paradigm. Part (b) outlines existing benchmarks and datasets with their data content (real or/and fictional, abbreviated as Fic.), experiment paradigm (with or without fine-tuning, abbreviated as FT), statistics of data (text/QA), evaluation tasks and applications in subsequent studies. The ‘+’ in the data column distinguishes between the unlearn set and the retain set. The meaning of abbreviations: “SAQA”: Short answer question & answer, “MCQA”: Multiple choice QA.
@@ 4.1.4. Existing Benchmarks and Datasets
A direct motivation for unlearning research comes from a number of works that aim to remove parts of the pretraining corpus (Yao et al., 2024a; Ashuach et al., 2025; Jang et al., 2023b; Wang et al., 2025c; Cha et al., 2024; Bărbulescu and Triantafillou, 2024; Stoehr et al., 2024; Pochinkov and Schoots, 2024; Tamirisa et al., 2024). Among these, the Pile dataset (Gao et al., 2020), which is commonly used in pretraining LLMs such as Pythia (Biderman et al., 2023), is one of the most frequently adopted. Google Research (2023) further introduced the Training Data Extraction Challenge (TDEC), a subset of 20,000 examples of The Pile that has been employed as an unlearn set in several studies. However, a major challenge is that the pretraining data for many state-of-the-art LLMs are not publicly available, and different models often use different corpora, significantly limiting the applicability of such datasets.
To address diverse research needs, numerous benchmarks and datasets have been developed, varying in motivation and application. Some focus on unlearning specific content, such as security information (Li et al., 2024; Lang et al., 2025), copyrighted material (Eldan and Russinovich, 2024; Shi et al., 2025), or private data (Liu et al., 2025b; Maini et al., 2024). Others emphasize knowledge connectivity or semantic diversity to enhance unlearning robustness (Qiu et al., 2024; Wu et al., 2025; Hong et al., 2024; Yang et al., 2025). Additionally, works such as (Rezaei et al., 2024; Veldanda et al., 2024) explore continuous learning–unlearning settings. We summarize the characteristics of existing benchmarks in Table 4.
Unlearning evaluations also frequently adapt benchmarks from related fields such as model editing and LLM safety. These include CounterFactual (Meng et al., 2022) (used by (Wang et al., 2025g; Guo et al., 2025; Patil et al., 2025)), PKU-SafeRLHF (Ji et al., 2023) (used by (Jia et al., 2025; Zhang et al., 2025a; Jia et al., 2024)), SQuAD (Rajpurkar et al., 2016) (used by (Pawelczyk et al., 2024; Rezaei et al., 2024)), and ZsRE (Yao et al., 2023) (used by (Wang et al., 2025b, g)).
@@ 4.2.1. Knowledge Memorization
In most cases, the ideal result of unlearning is expected to contain all of the retain set and none of the unlearn set. The first kind of metrics evaluates knowledge memorization, which examines whether certain data have been memorized in the model. Typically, the choice of knowledge memorization metric is tailored to the task format; for instance, accuracy is a direct measure for multiple choice questions. In general, metrics are categorized into two distinct classes according to their operational basis: those applied to the model’s final outputs and those applied to the model’s internal logits. We illustrate all the knowledge memorization metrics introduced in this section in Table 5.
@@
Output-based. The most direct approach is have the model complete the selected task and compare the output with the ground truth. Given that a model’s output may not perfectly align with ground-truth references, multiple metrics are employed to quantify the textual similarity between them. Verbatim matching represents the simplest and most computationally efficient approach, particularly suitable for short or categorical answers. For longer and more complex generations, studies such as (Thaker et al., 2024; Wang et al., 2025g) relax the exact match criterion to require the strict inclusion of specific keywords in the outputs. This adaptation demonstrates strong performance on benchmarks like TOFU (Maini et al., 2024), where questions are often centered on a unique, identifiable entity (e.g., an author’s name). Another way to relax is to check if there is a target token among the top-k tokens in the next token prediction process, and calculate metrics such as the top hit ratio (THR) (Qiu et al., 2024). BLEU (Papineni et al., 2002) and ROUGE (primarily ROUGE-N and ROUGE-L) (Lin, 2004) are established NLP metrics that focus on precision and recall of n-gram or longest common subsequence (LCS), respectively. However, they treat all words with equal weight. To prioritize key information, Xu et al. (2025) introduced the Entity Coverage Score (ECS), which extracts key entities using a model like deepseek-v3 and calculates similarity based solely on these entities.
@@
In addition to the model-free metrics, some methods introduce external models for a better evaluation beyond lexical overlap. Among these model-based methods, BertScore (Zhang et al., 2019) constitutes a major category, typically involving the conversion of text into embedding vectors and then the calculation of cosine similarity. Through this embedding transformation, the model can better handle semantic-level information, such as synonyms, negation words, and word order. For scenarios requiring more knowledge and understanding (such as recognizing that ”born in London” and ”born in the UK” are consistent), evaluation using NLI models can yield more accurate results. Furthermore, more universal evaluation methods include human evaluation or external LLM assessments. These methods are not only adaptable across diverse tasks, but can also comprehensively evaluate the wording and grammar of the outputs. However, they often function as black boxes and can be susceptible to biases inherent in human evaluators or proxy LLMs. Finally, several less prevalent metrics have also been applied in unlearning evaluations (Liu et al., 2024a; Gu et al., 2024), including METEOR (Banerjee and Lavie, 2005), MAUVE (Pillutla et al., 2021), and Rep3 (Welleck et al., 2019).
@@
$\rightarrow$ keyword (Thaker et al., 2024; Wang et al., 2025g), THR (Qiu et al., 2024), ES (Carlini et al., 2021)
(Ramakrishna et al., 2025; Cha et al., 2024; Thaker et al., 2024; Takashiro et al., 2024; Stoehr et al., 2024; Pochinkov and Schoots, 2024; Wang et al., 2025g; Tian et al., 2024; Wang et al., 2025d, 2024)
BLEU Commonly used in translation (precision). (Baluta et al., 2024; Wu et al., 2024; Hu et al., 2024; Liu et al., 2024a; Ji et al., 2024; Gu et al., 2024; Scholten et al., 2025)
@@
$\rightarrow$ ECS (Xu et al., 2025)
(Maini et al., 2024; Ramakrishna et al., 2025; Qiu et al., 2024; Joshi et al., 2024; Jiang et al., 2025b; Yuan et al., 2025b; Wei et al., 2025b; Scholten et al., 2025; Wang et al., 2025b; Shen et al., 2025; Zhang et al., 2025a; Cha et al., 2024; Hu et al., 2024; Chen and Yang, 2023; Ren et al., 2025; Huang et al., 2025; Liu et al., 2024a; Ji et al., 2024; Sanyal and Mandal, 2025; Wang et al., 2025e; Dou et al., 2025; Jia et al., 2025; Xu et al., 2025; Gu et al., 2024; Choi et al., 2025; Ma et al., 2025; Zhang et al., 2025e; Yuan et al., 2025a; Wang et al., 2025d; Mekala et al., 2025; Patil et al., 2025; Russinovich and Salem, 2025; Wei et al., 2025a; Liu et al., 2024b)
Model- based BertScore (Zhang et al., 2019) Semantic information (synonym, etc.).
@@
(Yuan et al., 2025b; Wei et al., 2025b; Wu et al., 2024; Liu et al., 2024a; Sanyal and Mandal, 2025; Xu et al., 2025; Russinovich and Salem, 2025; Wei et al., 2025a)
NLI Knowledge and understanding. (Liu et al., 2025b; Jiang et al., 2025b; Yuan et al., 2025b; Xu et al., 2025)
Human eval Adaptable across diverse tasks, comprehensive in multiple dimensions. (Liu et al., 2021; Xu et al., 2025)
LLM eval Similar to the notes of human eval. (Ramakrishna et al., 2025; Lang et al., 2025; Hu et al., 2024; Sanyal and Mandal, 2025; Wang et al., 2025e; Xu et al., 2025; Mekala et al., 2025; Liu et al., 2024b)
@@
e.g. $p(a|q)$ (Maini et al., 2024), perplexity
$\rightarrow$ KL divergence (Wang et al., 2025d), CI (Mekala et al., 2025)
(Eldan and Russinovich, 2024; Maini et al., 2024; Yao et al., 2024a; Doshi and Stickland, 2024; Ji et al., 2024; Wu et al., 2023; Wang et al., 2025d; Mekala et al., 2025; Patil et al., 2025)
Truth ratio (Maini et al., 2024) Detect under- and over-unlearning at the distribution level. (Zhang et al., 2025a; Wang et al., 2025b; Cha et al., 2024; Thaker et al., 2024; Huang et al., 2025; Liu et al., 2024a; Ji et al., 2024; Jia et al., 2025; Gu et al., 2024; Choi et al., 2025; Mekala et al., 2025; Jia et al., 2024; Wang et al., 2024; Patil et al., 2025)
@@
e.g. Exposure (Carlini et al., 2019), MRR (Lacroix et al., 2018)
(Ashuach et al., 2025; Wu et al., 2023; Baluta et al., 2024; Qiu et al., 2024)
@@
Logit-based. The autoregressive nature of Large Language Models (LLMs) involves computing a probability distribution for the next token conditioned on the preceding sequence. This next token probability can thus serve as an indicator of the model’s latent knowledge for a given prompt, where a higher probability signifies stronger retention of that specific token (Eldan and Russinovich, 2024; Maini et al., 2024). A representative indicator calculated on this basis is perplexity, which is adopted as a metric in several studies (Yao et al., 2024a; Doshi and Stickland, 2024; Ji et al., 2024; Wu et al., 2023), under the premise that more firmly memorized knowledge typically yields lower perplexity. A significant advancement is the Truth Ratio introduced by Maini et al. (2024), which quantifies the likelihood of a correct answer relative to incorrect alternatives for a given question. Theoretically, a model lacking specific knowledge should exhibit a negligible difference in probability between correct and incorrect answers. The efficacy of unlearning can be further statistically validated by applying tests like the KS-Test to compare the distribution of Truth Ratios between the unlearned model and an expected baseline. The truth ratio can effectively detect under- and over-unlearning at the distribution level.
A known limitation of direct probability usage is the extreme variance in conditional probabilities across tokens, which can adversely affect metric stability. A straightforward mitigation is to utilize token ranks instead of raw probabilities. Sorting tokens by their probability in descending order and using the rank as the score results offer a more uniform score distribution (Ashuach et al., 2025). This rank-based paradigm is also employed by several metrics adapted for unlearning evaluation (Wu et al., 2023; Baluta et al., 2024; Qiu et al., 2024). For instance, Exposure (Carlini et al., 2019), a key metric in memorization analysis, can be viewed as a rank-based variant of the Truth Ratio, substituting likelihood comparison with rank comparison. Similarly, the Mean Reciprocal Rank (MRR), prevalent in entity retrieval tasks (Lacroix et al., 2018), calculates the reciprocal average of the ranks of target tokens.

### v2

@@ 4.1.3. Experiment Paradigm
In the experiment, datasets can be broadly categorized into two classes based on whether fine-tuning is required. In the first category, the models perform unlearning without fine-tuning on the target dataset, which simplifies experimental setup. This includes the following scenarios. (1) The unlearn set is compiled directly from the model’s pretraining data, such as subsets derived from the Pile (gao2020pile). (2) The data are manually verified to be present in the model, as in RWKU (Jin2024RWKU), ConceptVectors (Hong2024Intrinsic), and RETURN (Liu2024Learning). However, verifying the presence of facts in LLMs remains challenging and may affect reliability. (3) For security purposes, the model is required to erase certain knowledge regardless of its original presence, exemplified by WMDP (Li2024WMDP) and UNCD (lang2025beyond). In the second category, models are first fine-tuned on the full dataset before a subset is unlearned. This is essential when datasets are fictionally synthetic, such as TOFU (Maini2024TOFU), EDU-RELAT (wu2024evaluating), and PISTOL (Qiu2024PISTOL), to ensure that the model acquires the target knowledge. Even for real-world corpora, fine-tuning helps to guarantee that the original model possesses knowledge of the unlearn set.
@@
TOFU (Maini2024TOFU) $0×\times2$0 QA SAQA (joshi2024towards; yuan2024closer; Scholten2024Probabilistic; wang2025erasing; shen2025lunar; zhang2025resolving; gao2024large; cha2024robust; ren2025general; thaker2024guardrail; Huang2024Offset; Liu2024Large; Ji2024Reversing; sanyal2025agents; takashiro2025answer; jia2024wagle; xu2025relearn; Gu2024MEOW; ma2025unveiling; wang2025effective; Mekala2024Alternate; Jia2024SOUL; Wang2024Rethinking; chen2025soft; Liu2024Revisiting; Bu2024Unlearning; Wang2024LLM; fan2025towards; fan2025simplicity; Wang2024RKLD; Zhang2024Negative)
WMDP (Li2024WMDP) Papers & passages 3,668 MCQA (che2025model; doshi2024does; wang2025erasing; farrell2024applying; ren2025general; thaker2024guardrail; Liu2024Large; Bhaila2024Soft; sanyal2025agents; kolbeinsson2025composable; jia2024wagle; hu2025falcon; huu2024effects; lucki2024adversarial; fan2025simplicity; Kadhe2024Split; deeb2025unlearning)
WHP (Eldan2023Whos) 3.1M tokens 300 Conti + 30 Cloze (Lynch2024Eight; Scholten2024Probabilistic; wang2025erasing; thaker2024guardrail; Ji2024Reversing; sanyal2025agents; jia2024wagle; Jia2024SOUL; chen2025soft)
MUSE (Shi2024MUSE) 4.4M+6.5M tokens Conti + SAQA (jiang2025holistic; wang2025erasing; zhang2025catastrophic; Yuan2024Robust; russinovich2025obliviate; Bu2024Unlearning; Wang2024LLM; fan2025simplicity; dong2024undial; zhang2025rule)
RWKU (Jin2024RWKU) 200 celebrities 3,268 cloze + 2,879 SAQA (takashiro2025answer; Yuan2024Robust; zhang2025rule)
CoTaEval (Wei2024Evaluating) 1K + 1K passages 1.5K Conti + 1K SAQA (russinovich2025obliviate)
KnowUnDo (Tian2024Forget) 2,649 QA SAQA (xu2025relearn)
PISTOL (Qiu2024PISTOL) 4 Graphs (50,95) 95 SAQA (shen2025lunar)
WPU (Liu2024Revisiting) 100 people’s Wiki 2,795 SAQA (Sinha2024UnStar)
LUME (ramakrishna2025lume) 1,387 documents 4,394 (Conti + SAQA) SEMEval-2025 Task 4
@@
ELUDe (choi2025opt-out) 15,651+90,954 QA MCQA + SAQA -
FaithUn (yang2025faithun) 664 QA 8,377 MCQA -
LLM Surgery (veldanda2024llm) 180K+1B tokens 24,800 MCQA -
Restor (rezaei2024restor) 3,000 passages 1,051 SAQA -
@@
UNCD (lang2025beyond) 2.9M+3.3M tokens 36K MCQA -
(b) Benchmarks and datasets in unlearning.
Benchmark Field Data Used by
CounterFact (meng2022locating) Editing 21,919 counterfactual records (Wang2024Large; guo2024mechanistic; Patil2025UPCORE)
PKU-SafeRLHF (ji2024pkusaferlhf) Safety QA pairs (265K w/ meta-labels, 166.8K w/ preference) (jia2024wagle; zhang2025resolving; Jia2024SOUL; Pan2025MultiObjective; chen2024wpn; chen2024machine; Liu2024Safer; Yao2024Large)
SQuAD (rajpurkar2016squad) Comprehension 100K+ QA pairs of reading and reasoning (Pawelczyk2024Context; rezaei2024restor)
ZsRE (levy2017zero) Relation extraction 30M+ positive examples, 2M+ negative examples (wang2025erasing; Wang2024Large)
(c) Benchmarks and datasets in relevant fields.
Table 3. Select a suitable benchmark. Part (a) organizes advantages of different tasks formats, data content and experiment paradigm. Part (b) outlines benchmarks and datasets with their data content (real or/and fictional (Fic.)), experiment paradigm (with or without fine-tuning (FT)), statistics of data (text/QA, ‘+’ distinguishes between the unlearn set and the retain set), evaluation tasks (“Conti”: Continuation, “SAQA”: Short answer QA, “MCQA”: Multiple choice QA) and applications in subsequent studies. Part (c) outlines benchmarks in relevant fields with brief descriptions of their data.
@@ 4.1.4. Existing Benchmarks and Datasets
A direct motivation for unlearning research comes from a number of works that aim to remove parts of the pretraining corpus (yao2024machine; Ashuach2024REVS; Wang2024Selective; cha2024robust; Stoehr2024Localizing; pochinkov2024dissecting; Tamirisa2024Robust; belrose2023leace). Among these, the Pile dataset (gao2020pile), which is commonly used in pretraining LLMs such as Pythia (biderman2023pythia), is one of the most frequently adopted. lm_extraction_benchmark further introduced the Training Data Extraction Challenge (TDEC), a subset of 20,000 examples of The Pile that has been employed as an unlearn set in several studies (jang2022knowledge; Tang2024Learn; Lee2024Protecting; barbulescu2024each; Gao2024Ethos; dong2024undial; Wang2024Selective; kassem2023preserving). However, a major challenge is that the pretraining data for many state-of-the-art LLMs are not publicly available, and different models often use different corpora, significantly limiting the applicability of such datasets.
To address diverse research needs, numerous benchmarks and datasets have been developed, varying in motivation and application. Some focus on unlearning specific content, such as security information (Li2024WMDP; lang2025beyond), copyrighted material (Eldan2023Whos; Shi2024MUSE), or private data (Liu2024Learning; Maini2024TOFU). Others emphasize knowledge connectivity or semantic diversity to enhance unlearning robustness (Qiu2024PISTOL; wu2024evaluating; Hong2024Intrinsic; yang2025faithun). Additionally, works such as (rezaei2024restor; veldanda2024llm) explore continuous learning–unlearning settings. Unlearning evaluations also frequently adapt benchmarks from related fields such as model editing and LLM safety. We summarize the characteristics of existing benchmarks in Table 3.
@@ 4.2.1. Knowledge Memorization
In most cases, the ideal result of unlearning is expected to contain all of the retain set and none of the unlearn set. The first kind of metrics evaluates knowledge memorization, which examines whether certain data have been memorized in the model. Typically, the choice of knowledge memorization metric is tailored to the task format; for instance, accuracy is a direct measure for multiple choice questions. In general, metrics are categorized into three distinct classes according to their operational basis: those applied to the model’s final outputs, those applied to the model’s internal logits, and membership inference attack (MIA). We summarize the knowledge memorization metrics introduced in this section in Table 4.
Output-based
@@
The most direct approach is have the model complete the selected task and compare the output with the ground truth. Given that a model’s output may not perfectly align with ground-truth references, multiple metrics are employed to quantify the textual similarity between them. Verbatim matching represents the simplest and most computationally efficient approach, particularly suitable for short or categorical answers. For longer and more complex generations, studies such as (thaker2024guardrail; Wang2024Large) relax the exact match criterion to require the strict inclusion of specific keywords in the outputs. This adaptation demonstrates strong performance on benchmarks like TOFU (Maini2024TOFU), where questions are often centered on a unique, identifiable entity (e.g., an author’s name). Another way to relax is to prompt the first $n$ tokens and check the $n+1$ token at each time, including memorization accuracy (MA) (tirumala2022memorization) and extraction strength (ES) (carlini2021extracting). BLEU (papineni2002bleu) and ROUGE (primarily ROUGE-N and ROUGE-L) (lin2004rouge) are established NLP metrics that focus on precision and recall of n-gram or longest common subsequence (LCS), respectively. Beyond a single score, Extraction Likelihood (EL) (jang2022knowledge) measures extraction risk as the average ROUGE score across varying prefix lengths. However, both ROUGE and EL treat all words with equal weight. To prioritize key information, xu2025relearn introduced the Entity Coverage Score (ECS), which extracts key entities by an LLM and calculates similarity based solely on these entities.
@@
In addition to the model-free metrics, some methods introduce external models for a better evaluation beyond lexical overlap. Among these model-based methods, BertScore (zhang2019bertscore) constitutes a major category, typically involving the conversion of text into embedding vectors and then the calculation of cosine similarity. Through this embedding transformation, the model can better handle semantic-level information, such as synonyms, negation words, and word order. For scenarios requiring more knowledge and understanding (such as recognizing that ”born in London” and ”born in the UK” are consistent), evaluation using customized models, such as Natural Language Inference (NLI) models (bos2005recognising) or BLEURT (sellam2020bleurt), can yield more accurate results. Furthermore, more universal evaluation methods include human evaluation or external LLM assessments. These methods are not only adaptable across diverse tasks, but can also comprehensively evaluate the wording and grammar of the outputs. However, they often function as black boxes and can be susceptible to biases inherent in human evaluators or proxy LLMs. Finally, several less prevalent metrics have also been applied in unlearning evaluations (Liu2024Large; Gu2024MEOW), including METEOR (banerjee2005meteor), MAUVE (pillutla2021mauve), and Rep3 (welleck2019neural).
Logit-based
The autoregressive nature of Large Language Models (LLMs) involves computing a probability distribution for the next token conditioned on the preceding sequence. This next token probability can thus serve as an indicator of the model’s latent knowledge for a given prompt, where a higher probability signifies stronger retention of that specific token (Eldan2023Whos; Maini2024TOFU). A representative indicator calculated on this basis is perplexity, which is adopted as a metric in several studies (yao2024machine; doshi2024does; Ji2024Reversing; Wu2023DEPN; choi2024cross), under the premise that more firmly memorized knowledge typically yields lower perplexity. Several work has calculated comprehensive indicators based on the next token probability (wang2025effective; Mekala2024Alternate; Lee2024Protecting). A significant advancement is the Truth Ratio introduced by Maini2024TOFU, which quantifies the likelihood of a correct answer relative to incorrect alternatives for a given question. Theoretically, a model lacking specific knowledge should exhibit a negligible difference in probability between correct and incorrect answers. The efficacy of unlearning can be further statistically validated by applying tests like the KS-Test to compare the distribution of Truth Ratios between the unlearned model and an expected baseline. The truth ratio can effectively detect under- and over-unlearning at the distribution level.
A known limitation of direct probability usage is the extreme variance in conditional probabilities across tokens, which can adversely affect metric stability. A straightforward mitigation is to utilize token ranks instead of raw probabilities. Sorting tokens by their probability in descending order and using the rank as the score results offer a more uniform score distribution (Ashuach2024REVS). This rank-based paradigm is also employed by several metrics adapted for unlearning evaluation (Qiu2024PISTOL; choi2024cross). For instance, Exposure (carlini2019secret), a key metric in memorization analysis, can be viewed as a rank-based variant of the Truth Ratio, substituting likelihood comparison with rank comparison. Similarly, the Mean Reciprocal Rank (MRR), prevalent in entity retrieval tasks (lacroix2018canonical), calculates the reciprocal average of the ranks of target tokens.
Membership inference attack (MIA)
In addition to the metrics above, some articles conduct MIA for evaluation (Jin2024RWKU; Shi2024MUSE; ramakrishna2025lume; Jia2024SOUL). MIA is a privacy attack that determines whether specific data samples are part of a model’s training set (shokri2017membership), including LOSS (yeom2018privacy), Zlib Entropy (carlini2021extracting), Min-K% Prob (shi2023detecting) and Min-K%++ Prob (zhang2024min). However, although widely used in traditional unlearning evaluations, MIA is considered unsuitable for the LLM context, as it typically requires training numerous shadow models, which is both data-prohibitive and computationally intractable for LLMs (Liu2024Large; chen2024machine; Yao2024Large).
@@
$\rightarrow$ keyword (thaker2024guardrail; Wang2024Large), ES (carlini2021extracting), MA (tirumala2022memorization)
(ramakrishna2025lume; cha2024robust; thaker2024guardrail; takashiro2025answer; Stoehr2024Localizing; pochinkov2024dissecting; Wang2024Large; Tian2024Forget; wang2025effective; Wang2024Rethinking; jang2022knowledge; Tang2024Learn; Lee2024Protecting; choi2024cross; barbulescu2024each; dong2024undial; Wang2024Selective)
BLEU Commonly used in translation (precision).
$\rightarrow$ Scare-BLEU (post2018call)
(baluta2024unlearning; wu2024codeunlearn; hu2024separate; Liu2024Large; Ji2024Reversing; Gu2024MEOW; Scholten2024Probabilistic; Wang2024LLM; Tang2024Learn; Yao2024Large; Wang2023KGA)
@@
$\rightarrow$ EL (jang2022knowledge), ECS (xu2025relearn)
(Maini2024TOFU; ramakrishna2025lume; Qiu2024PISTOL; joshi2024towards; jiang2025holistic; yuan2024closer; wei2024underestimated; Scholten2024Probabilistic; wang2025erasing; shen2025lunar; zhang2025resolving; cha2024robust; hu2024separate; Chen2023Unlearn; ren2025general; Huang2024Offset; Liu2024Large; Ji2024Reversing; sanyal2025agents; wang2024when; Dou2024Avoiding; jia2024wagle; xu2025relearn; Gu2024MEOW; choi2025opt-out; ma2025unveiling; zhang2025catastrophic; Yuan2024Robust; wang2025effective; Mekala2024Alternate; Patil2025UPCORE; russinovich2025obliviate; Wei2024Evaluating; Liu2024Revisiting; Bu2024Unlearning; Sinha2024UnStar; Wang2024LLM; fan2025simplicity; jang2022knowledge; Tang2024Learn; Lee2024Protecting; Wang2024RKLD; Zhang2024Negative; Gao2024Ethos; dong2024undial; Wang2024Selective; zhang2025rule)
Model- based BertScore (zhang2019bertscore) Semantic information (synonym, etc.).
@@
(yuan2024closer; wei2024underestimated; wu2024codeunlearn; Liu2024Large; sanyal2025agents; xu2025relearn; russinovich2025obliviate; Wei2024Evaluating; Tang2024Learn; Yao2024Large)
Customized model eval Knowledge and understanding.
e.g. NLI, BLEURT (sellam2020bleurt)
(Liu2024Learning; jiang2025holistic; yuan2024closer; xu2025relearn; chen2024machine; Liu2024Safer; Yao2024Large)
Human eval Adaptable across diverse tasks, comprehensive in multiple dimensions. (liu2021dexperts; xu2025relearn; Tang2024Learn; Lu2022QUARK)
LLM eval Similar to the notes of human eval. (ramakrishna2025lume; lang2025beyond; hu2024separate; sanyal2025agents; wang2024when; xu2025relearn; Mekala2024Alternate; Liu2024Revisiting; Pan2025MultiObjective; Sinha2024UnStar; Tang2024Learn; Liu2024Safer; Lu2022QUARK; zhang2025rule)
@@
e.g. $p(a|q)$ (Maini2024TOFU), perplexity
$\rightarrow$ KL divergence (wang2025effective), CI (Mekala2024Alternate), RMA (Lee2024Protecting)
(Eldan2023Whos; Maini2024TOFU; yao2024machine; doshi2024does; Ji2024Reversing; Wu2023DEPN; wang2025effective; Mekala2024Alternate; Patil2025UPCORE; fan2025simplicity; Lee2024Protecting; choi2024cross; Wang2024RKLD; Zhang2024Negative; Wang2023KGA; Lu2022QUARK)
Truth ratio (Maini2024TOFU) Detect under- and over-unlearning at the distribution level. (zhang2025resolving; wang2025erasing; cha2024robust; thaker2024guardrail; Huang2024Offset; Liu2024Large; Ji2024Reversing; jia2024wagle; Gu2024MEOW; choi2025opt-out; Mekala2024Alternate; Jia2024SOUL; Wang2024Rethinking; Patil2025UPCORE; Bu2024Unlearning; Wang2024LLM; fan2025simplicity; Wang2024RKLD; Zhang2024Negative)
@@
e.g. Exposure (carlini2019secret), MRR (lacroix2018canonical), THR (Qiu2024PISTOL), PA (choi2024cross)
(Ashuach2024REVS; Wu2023DEPN; baluta2024unlearning; Qiu2024PISTOL; choi2024cross; Gu2024SecondOrder)